# The evolution of the galaxy stellar mass function and star formation rates in the COLIBRE simulations from redshift 17 to 0

### v1

@@ 3.7 The abundance of massive quiescent galaxies at high redshift
Fig. 11 shows the redshift evolution of the number density of quiescent galaxies at $1.75<z<9.25$. Unlike in Figs. 9 and 10, here we define quenched galaxies as those with $\mathrm{sSFR}<10^{-10}~\mathrm{yr}^{-1}$, to facilitate a fair comparison with the observational data (discussed later). We present the number densities of quiescent galaxies in colibre for several stellar mass bins commonly used in observational studies: $M_{*}>10^{10}~\mathrm{M_{\odot}}$ (top panel), $M_{*}>10^{10.5}~\mathrm{M_{\odot}}$ (middle panel), and $M_{*}>10^{11}~\mathrm{M_{\odot}}$ (bottom panel). The shaded regions indicate Poisson uncertainties in the simulation predictions.
Unlike in previous figures, we show only the L400m7 and L200m6 simulations, as the volume of the largest m5 simulation ($50^{3}~\mathrm{cMpc}^{3}$) is too small for a meaningful comparison (see Fig. 15 in Appendix A). Following $\S$2.3, the number densities predicted by the simulations are computed using galaxy stellar masses with a lognormal scatter (equation 1) to account for Eddington bias. The effect of Eddington bias on the number density of quiescent galaxies is discussed in Appendix D, where we show that adding the scatter improves the agreement with observations (see Fig. 21).
For comparison, in Fig. 11 we include JWST data from Valentino et al. (2023), Nanayakkara et al. (2024), Russell et al. (2024), Weibel et al. (2025b), de Graaff et al. (2025), Baker et al. (2025a), and Zhang et al. (2025). Briefly,
@@
  • Baker et al. (2025a) analyzed a sample of 745 massive quiescent galaxies over the redshift range $2<z<7$, using JWST/NIRCam imaging from a compilation of public JWST fields covering a combined area of over $800$ arcmin2. Photometric redshifts were derived using eazy, and quiescent galaxies were identified based on an sSFR threshold of $\mathrm{sSFR}<0.2/t_{\rm age}(z)$. For our comparison, we consider only the subset of their sample with stellar masses $M_{*}>10^{10}~\mathrm{M_{\odot}}$.
@@
Figure 11: Evolution of the comoving number density of massive quiescent galaxies in the colibre L400m7 and L200m6 simulations. Quiescent galaxies are defined as those with $\mathrm{sSFR}<10^{-10}~\mathrm{yr}^{-1}$, mimicking observational selection criteria at high redshifts. We show the number densities of quiescent galaxies with stellar masses of $M_{*}>10^{10}~\mathrm{M_{\odot}}$ (top panel), $M_{*}>10^{10.5}~\mathrm{M_{\odot}}$ (middle panel), and $M_{*}>10^{11}~\mathrm{M_{\odot}}$ (bottom panel). The shaded regions indicate the Poisson uncertainty on the simulation predictions. For comparison, we include JWST data with similar $M_{*}$ and sSFR thresholds from Valentino et al. (2023); Weibel et al. (2025b); Nanayakkara et al. (2024); Russell et al. (2024), de Graaff et al. (2025), Baker et al. (2025a), and Zhang et al. (2025). All JWST data up to $z\approx 7$ are consistent with at least one of the colibre simulations, although the m6 model systematically predicts a lower number density of quiescent galaxies than m7 at fixed redshift.
@@
We find that the number density of massive quiescent galaxies, $n_{\rm q}$, predicted by the colibre simulations increases steeply with decreasing redshift, rising by more than four orders of magnitude between $z=9$ and $z=2$ in all three stellar mass bins. Furthermore, comparing the different panels at fixed redshift, we see that $n_{\rm q}$ decreases as the stellar mass threshold increases from $M_{*}=10^{10}~\mathrm{M_{\odot}}$ to $10^{11}~\mathrm{M_{\odot}}$. Thanks to the relatively large volumes of the m7 and m6 simulations, quiescent galaxies with $M_{*}>10^{10}~\mathrm{M_{\odot}}$ are sampled out to $z=9$ and $z=7.5$ in the m7 and m6 simulations, respectively. For the most massive quiescent galaxies ($M_{*}>10^{11}~\mathrm{M_{\odot}}$), the m7 (m6) simulation predicts a non-zero number density at $z\leq 7.5$ ($z\leq 5$).
@@
When comparing the m6 and m7 simulations at fixed redshift and for the same stellar mass bin, we find that the predicted values of $n_{\rm q}$ are systematically higher in the m7 simulation, with differences ranging from $\approx 0.2$ to $\approx 1$ dex. As shown in Appendix A, these discrepancies cannot be due to the eight times larger volume of the m7 simulation. Instead, they arise from weaker, more gradual quenching at high redshifts in the higher-resolution colibre simulations – a trend already observed in the evolution of the galaxy quenched fraction (see Fig. 9).
@@
All JWST measurements are reproduced by at least one of the two colibre simulations (m6 or m7). In the $M_{*}>10^{10}~\mathrm{M_{\odot}}$ mass bin, the m6 and m7 simulations both predict $n_{\rm q}\sim 10^{-6.5}~\mathrm{cMpc}^{-3}$ at $z\approx 7$, which lies within $\approx 1\sigma$ of the measurements by Weibel et al. (2025b) and Baker et al. (2025a). The m6 simulation is also consistent with Baker et al. (2025a) at all lower redshifts probed by the observations ($2<z<6$), as well as with the measurement from Nanayakkara et al. (2024) at $z\approx 3.5$. In contrast, the m7 model overpredicts the observed number densities from Baker et al. (2025a) and Nanayakkara et al. (2024) by $\approx 0.25$ dex to $\approx 1$ dex at these redshifts. In the $M_{*}>10^{10.5}~\mathrm{M_{\odot}}$ bin, the m7 and m6 simulations predict quiescent galaxy number densities of $n_{\rm q}\sim 10^{-6}~\mathrm{cMpc}^{-3}$ at $z\approx 6$ and $z\approx 5$, respectively, rising in both cases to $\sim 10^{-3.5}~\mathrm{cMpc}^{-3}$ by $z\approx 2$. The high-redshift measurements reported by Valentino et al. (2023) and Russell et al. (2024) at $z\approx 6$, as well as by Zhang et al. (2025) at $z\approx 4.5$, are reproduced by the m7 simulation, while the m6 simulation shows good agreement with the lower-redshift bins ($z\lesssim 4$) reported in these studies. For the highest mass bin, the m7 (m6) simulation predicts a number density of $\approx 1.5\times 10^{-6}~\mathrm{cMpc}^{-3}$ ($\approx 10^{-7}~\mathrm{cMpc}^{-3}$) at $z\approx 4.9$, placing it within $\approx 1.5\sigma$ ($\approx 3\sigma$) of the single galaxy-based estimate by de Graaff et al. (2025).
@@
We note that earlier galaxy formation simulations and semi-analytic models that predict reasonably realistic low-redshift galaxy populations typically strongly underpredict the number density of massive quiescent galaxies observed by JWST at $z\gtrsim 5$ (see, e.g., fig. 6 in Weibel et al. 2025b and the section ‘Theoretical predictions’ in de Graaff et al. 2025). The improved performance of colibre in this metric may be attributed to several factors. First, unlike most earlier galaxy simulations, colibre does not implement an effective equation of state (EoS) to model the ISM. The absence of an EoS results in denser and cooler gas within the ISM, which can enable more rapid BH accretion. Consequently, the more efficiently growing SMBHs may trigger stronger AGN feedback, facilitating the quenching of massive galaxies at earlier times. Second, colibre does not impose a cap on SMBH accretion at the Eddington rate. This approach also allows faster SMBH growth and potentially stronger quenching, particularly at high redshifts where super-Eddington accretion events are expected to be more common (e.g. Huško et al., 2025b). A detailed investigation of the relevance of these factors will be conducted in future work. Finally, as explained in $\S$2.3, galaxy stellar masses in this work include a lognormal scatter with a standard deviation given by equation (1) to account for Eddington bias, which is present in observational data due to non-negligible uncertainties in the inferred stellar masses. As we show in Fig. 21 in Appendix D, the number density of massive quiescent galaxies is particularly sensitive to this scatter, which on average boosts the value of $n_{\rm q}$ at fixed redshift and stellar mass bin. This effect may not have been fully accounted for in some previous works comparing simulations with observations, potentially biasing simulation-based estimates of $n_{\rm q}$ towards lower values.
@@ 3.8 Comparison of models with thermal and hybrid AGN feedback
Thus far, we have discussed exclusively the colibre simulations with thermal AGN feedback. However, the colibre suite also includes models with hybrid (thermal + kinetic jet) AGN feedback. In this section, we make use of these additional simulations to investigate how differences in AGN feedback prescriptions impact the key galaxy properties examined in this study: stellar masses, SFRs, and quenched fractions. We emphasize that while both the thermal and hybrid models were calibrated to reproduce the $z=0$ GSMF from Driver et al. (2022) and the $z\approx 0$ stellar mass – size relation from Hardwick et al. (2022), the properties of galaxies other than their $z=0$ stellar masses and half-mass radii, should not necessarily be expected to match, especially at higher redshifts where no observational constraints were applied to tune the simulations. Additionally, although the primary difference between the thermal and hybrid models lies in the AGN feedback prescription, we note that one of the SN feedback parameters also differs slightly. Namely, at all resolutions, the normalization of the gas density-dependent heating temperature $\Delta T_{\rm SN}(\rho_{\rm gas})\propto\rho_{\rm gas}^{2/3}$ is increased in the hybrid model, resulting in lower heating temperatures at fixed gas density and therefore weaker SN feedback. This adjustment was introduced to counterbalance the stronger AGN feedback in the hybrid model in intermediate-mass galaxies (see Huško et al. 2025a for details).
@@
  • At $z=0.1$, the GSMFs of the thermal and hybrid models become indistinguishable once again, while the trend in quenched fractions reverses. The hybrid model now exhibits higher quenched fractions, reaching $\approx 35$ per cent at $10^{9}\lesssim M_{*}\lesssim 10^{10}~\mathrm{M_{\odot}}$, compared to $\approx 20$ per cent in the thermal model. Additionally, the normalization of the SFMS is lower in the hybrid model by about $0.1$ dex. At higher stellar masses ($M_{*}\gtrsim 10^{11}~\mathrm{M_{\odot}}$), the quenched fractions in both models converge, with the hybrid model predicting marginally higher values. At even greater masses ($M_{*}\gtrsim 10^{11.5}~\mathrm{M_{\odot}}$), the quenched fractions in both models decline to around 50 per cent, suggesting that neither thermal energy injection nor kinetic jets can fully offset the strong cooling flows in the most massive galaxies. However, we note that the quenched fractions at the high-mass end are particularly sensitive to uncertainties in stellar mass due to Eddington bias, which may artificially elevate their values (see Appendix D).
The hybrid model exhibits lower quenched fractions at high redshifts due to less aggressive AGN feedback, resulting from the slower BH growth (see Huško et al., in preparation, for details). Over time, these lower quenched fractions, combined with slightly weaker SN feedback, allow more stellar mass to form – resulting in a marginally elevated GSMF in the hybrid model at $M_{*}\gtrsim 10^{9.5}~\mathrm{M_{\odot}}$ by $z=2$. By $z=0.1$, the GSMFs in both models converge, as they were calibrated to match the same observational constraints. Mathematically, the lowz-z quenching in the hybrid model is expected to be stronger than in the thermal model, to ensure that the same amount of stellar mass is formed by $z=0$. Physically, this stronger quenching arises from the hybrid model’s more efficient AGN feedback, which at low redshifts is dominated by the kinetic jet mode, which experiences less radiative losses compared to thermal AGN feedback.
@@
Both models form similar amounts of stellar mass at $z\gtrsim 7$. Between $z=7$ and $z=2$, the CSMDs in the two models diverge, with the differences peaking at $z\approx 2$, where the hybrid model has formed $\approx 25$ per cent more stellar mass. From $z=2$ to 0, the CSMDs converge, with the differences approaching zero by $z\approx 0$, as both models were calibrated to reproduce the observed $z=0$ GSMF.

### v2

@@ 3.7 The abundance of massive quiescent galaxies at high redshift
Fig. 11 shows the redshift evolution of the number density of quiescent galaxies at $1.75<z<9.25$. Unlike in Figs. 9 and 10, here we define quenched galaxies as those with $\mathrm{sSFR}<10^{-10}~\mathrm{yr}^{-1}$, to facilitate a fair comparison with the observational data (discussed later). We present the number densities of quiescent galaxies in colibre for several stellar mass bins commonly used in observational studies: $M_{*}>10^{10}~\mathrm{M_{\odot}}$ (top panel), $M_{*}>10^{10.5}~\mathrm{M_{\odot}}$ (middle panel), and $M_{*}>10^{11}~\mathrm{M_{\odot}}$ (bottom panel). The shaded regions indicate Poisson uncertainties in the simulation predictions. We show the predictions from the largest colibre simulations at each resolution: L400m7, L200m6, and L100m5 (colours). Unlike in previous figures, we do not switch from the L100m5 simulation to L050m5 below $z=4.6$, as the $50^{3}~\mathrm{cMpc}^{3}$ volume is too small for a meaningful comparison (see Fig. 15 in Appendix A).
Following $\S$2.3, the number densities predicted by the simulations are computed using galaxy stellar masses with a lognormal scatter (equation 1) to account for Eddington bias. The effect of Eddington bias on the number density of quiescent galaxies is discussed in Appendix D, where we show that adding the scatter improves the agreement with observations (see Fig. 21). At very high redshift, the number of quiescent objects may fluctuate between zero and a few because of small-number statistics. In such cases, we use upward-pointing arrows (instead of plotting a continuous curve) to indicate redshift bins where the number density of quiescent galaxies is greater than zero but both adjacent bins are zero. The arrows are coloured consistently with the curves corresponding to the different resolutions, with their tips marking the corresponding non-zero number density values.
For comparison, in Fig. 11 we include JWST data from Valentino et al. (2023), Nanayakkara et al. (2024), Russell et al. (2025), Weibel et al. (2025b), de Graaff et al. (2025), Baker et al. (2025b), and Zhang et al. (2025). Briefly,
@@
  • Baker et al. (2025b) analyzed a sample of 745 massive quiescent galaxies over the redshift range $2<z<7$, using JWST/NIRCam imaging from a compilation of public JWST fields covering a combined area of over $800$ arcmin2. Photometric redshifts were derived using eazy, and quiescent galaxies were identified based on an sSFR threshold of $\mathrm{sSFR}<0.2/t_{\rm age}(z)$. For our comparison, we consider only the subset of their sample with stellar masses $M_{*}>10^{10}~\mathrm{M_{\odot}}$.
@@
Figure 11: Evolution of the comoving number density of massive quiescent galaxies in the colibre L400m7, L200m6, and L100m5 simulations (coloured curves). Quiescent galaxies are defined as those with $\mathrm{sSFR}<10^{-10}~\mathrm{yr}^{-1}$, mimicking observational selection criteria at high redshift. We show the number densities of quiescent galaxies with stellar masses $M_{*}>10^{10}~\mathrm{M_{\odot}}$ (top panel), $M_{*}>10^{10.5}~\mathrm{M_{\odot}}$ (middle panel), and $M_{*}>10^{11}~\mathrm{M_{\odot}}$ (bottom panel). The shaded regions indicate the Poisson uncertainties on the simulation predictions. The upward-pointing arrows denote non-zero number density values in the regime of small-number statistics at high redshift and are coloured consistently with the corresponding curves (see text for details). For comparison, we include JWST data with similar $M_{*}$ and sSFR thresholds from Valentino et al. (2023); Weibel et al. (2025b); Nanayakkara et al. (2024); Russell et al. (2025), de Graaff et al. (2025), Baker et al. (2025b), and Zhang et al. (2025). All JWST data up to $z\approx 7$ are consistent with at least one of the colibre simulations, although the m6 model systematically predicts a lower number density of quiescent galaxies than m7 at fixed redshift.
@@
We find that the number density of massive quiescent galaxies, $n_{\rm q}$, predicted by the colibre simulations increases steeply with decreasing redshift, rising by more than four orders of magnitude between $z=9$ and $z=2$ in all three stellar mass bins. Furthermore, comparing the different panels at fixed redshift, we see that $n_{\rm q}$ decreases as the stellar mass threshold increases from $M_{*}=10^{10}~\mathrm{M_{\odot}}$ to $10^{11}~\mathrm{M_{\odot}}$. Thanks to the relatively large volumes of the m7 and m6 simulations, quiescent galaxies with $M_{*}>10^{10}~\mathrm{M_{\odot}}$ are sampled out to $z=9$ and $z=7.5$ in the m7 and m6 simulations, respectively. For the most massive quiescent galaxies ($M_{*}>10^{11}~\mathrm{M_{\odot}}$), the m7 (m6) simulation predicts a non-zero number density at $z\leq 7.5$ ($z\leq 6$). The predictions from the m5 simulation are present only over a very narrow redshift interval, $5<z<6$, and only for the lowest stellar mass bin ($M_{*}>10^{10}~\mathrm{M_{\odot}}$). This is expected: at $z\lesssim 5$ the simulation is not yet available, while at higher redshifts the $(100~\mathrm{cMpc})^{3}$ volume of the m5 simulation is too small to yield robust predictions, particularly in the higher stellar mass bins.
@@
When comparing the m6 and m7 simulations at fixed redshift and for the same stellar mass bin, we find that the predicted values of $n_{\rm q}$ are systematically higher in the m7 simulation, with differences ranging from $\approx 0.2$ to $\approx 1$ dex [^12] Unlike the m6 and m7 models, the m5 and m6 simulations formally converge for the lowest stellar mass bin in the narrow redshift range where the predictions of the L100m5 simulation are available ($5<z<6$). However, given the tendency of the $100^{3}~\mathrm{cMpc}^{3}$ volume to produce higher $n_{\rm q}$ than larger volumes at these redshifts (see Fig. 15), the predictions of the m5 model in a larger volume would likely lie slightly below those of L200m6.. As shown in Appendix A, these discrepancies cannot be due to the eight times larger volume of the m7 simulation. Instead, they arise from weaker, more gradual quenching at high redshifts in the higher-resolution colibre simulations – a trend already observed in the evolution of the galaxy quenched fraction (see Fig. 9). As discussed in $\S$3.6, a likely explanation for this behaviour is stronger radiative cooling losses at higher resolution, combined with less bursty AGN feedback, leading to overall weaker AGN feedback and consequently weaker quenching of massive galaxies.
@@
All JWST measurements are reproduced by at least one of the colibre simulations. In the $M_{*}>10^{10}~\mathrm{M_{\odot}}$ mass bin, the m6 and m7 simulations both predict $n_{\rm q}\sim 10^{-6.5}~\mathrm{cMpc}^{-3}$ at $z\approx 7$, which lies within $\approx 1\sigma$ of the measurements by Weibel et al. (2025b) and Baker et al. (2025b). The m6 simulation is also consistent with Baker et al. (2025b) at all lower redshifts probed by the observations ($2<z<6$), as well as with the measurement of Nanayakkara et al. (2024) at $z\approx 3.5$. The m5 model, where available, is likewise consistent with Baker et al. (2025b). In contrast, the m7 model overpredicts the observationally inferred number densities from Baker et al. (2025b) and Nanayakkara et al. (2024) by $\approx 0.25$ dex to $\approx 1$ dex at these redshifts. In the $M_{*}>10^{10.5}~\mathrm{M_{\odot}}$ bin, the m7 and m6 simulations predict quiescent galaxy number densities of $n_{\rm q}\sim 10^{-6}~\mathrm{cMpc}^{-3}$ at $z\approx 6$ and $z\approx 5$, respectively, rising in both cases to $\sim 10^{-3.5}~\mathrm{cMpc}^{-3}$ by $z\approx 2$. The high-redshift measurements reported by Valentino et al. (2023) and Russell et al. (2025) at $z\approx 6$, as well as by Zhang et al. (2025) at $z\approx 4.5$, are reproduced by the m7 simulation, while the m6 simulation shows good agreement with the lower-redshift bins ($z\lesssim 4$) reported in these studies. For the highest mass bin, the m7 (m6) simulation predicts a number density of $\approx 1.5\times 10^{-6}~\mathrm{cMpc}^{-3}$ ($\approx 10^{-7}~\mathrm{cMpc}^{-3}$) at $z\approx 4.9$, placing it within $\approx 1.5\sigma$ ($\approx 3\sigma$) of the single galaxy-based estimate by de Graaff et al. (2025).
@@
We note that earlier galaxy formation simulations and semi-analytic models that predict reasonably realistic low-redshift galaxy populations typically strongly underpredict the number density of massive quiescent galaxies observed by JWST at $z\gtrsim 5$ (see, e.g., fig. 6 in Weibel et al. 2025b and the section ‘Theoretical predictions’ in de Graaff et al. 2025). The improved performance of colibre in this metric may be attributed to several factors. First, unlike most earlier galaxy simulations, colibre does not implement an effective equation of state (EoS) to model the ISM. The absence of an EoS results in denser and cooler gas within the ISM, which can enable more rapid BH accretion. Consequently, the more efficiently growing SMBHs may trigger stronger AGN feedback, facilitating the quenching of massive galaxies at earlier times. Second, colibre does not impose a cap on SMBH accretion at the Eddington rate. This approach also allows faster SMBH growth and potentially stronger quenching, particularly at high redshifts where super-Eddington accretion events are expected to be more common (e.g. Huško et al., 2025). A detailed investigation of the relevance of these factors will be conducted in future work.
Furthermore, as explained in $\S$2.3, galaxy stellar masses in this work include a lognormal scatter with a standard deviation given by equation (1) to account for Eddington bias, which is present in observational data due to non-negligible uncertainties in the inferred stellar masses. As we show in Fig. 21 in Appendix D, the number density of massive quiescent galaxies is particularly sensitive to this scatter, which on average boosts the value of $n_{\rm q}$ at fixed redshift and stellar mass bin. This effect may not have been fully accounted for in some previous works comparing simulations with observations, potentially biasing simulation-based estimates of $n_{\rm q}$ towards lower values.
Lastly, we note that many earlier galaxy formation simulations were conducted in cosmological volumes of $\sim 100^{3}~\mathrm{cMpc}^{3}$, which strongly limits the non-zero number density of quenched galaxies that can be predicted at high redshift. For example, Lovell et al. (2023) showed that the highest redshift at which a non-zero abundance of massive quenched systems is found in the eagle simulations increases from $z\approx 5$ in the original eagle volume of $100^{3}~\mathrm{cMpc}^{3}$ to $z\approx 8$ in the Flares simulations, which are zoom-in simulations selected from a parent volume of $3.2^{3}~\mathrm{Gpc}^{3}$ and employing the same eagle galaxy formation model.
We emphasize that the number density of quiescent galaxies presented in this work is computed using an $\mathrm{sSFR}<10^{-10}~\mathrm{yr}^{-1}$ threshold to define quiescent galaxies. In all but one of the observational studies that we show for comparison in Fig. 11, galaxies were selected using (or satisfy) a similar sSFR threshold (see precise definitions above) to maximize consistency in the comparison and minimize potential bias in our conclusions. Furthermore, very recently, our results were confirmed by Chandro-Gómez et al. (2025), who conducted a detailed study of the properties of massive quenched galaxies in colibre at high redshift, finding generally good agreement with JWST data while adopting $\mathrm{sSFR}<0.2/t_{\rm age}(z)$ to define quiescent systems.
@@ 3.8 Comparison of models with thermal and hybrid AGN feedback
Thus far, we have discussed exclusively the colibre simulations with thermal AGN feedback. However, the colibre suite also includes models with hybrid (thermal + kinetic jet) AGN feedback. In this section, we make use of these additional simulations to investigate how differences in AGN feedback prescriptions impact the key galaxy properties examined in this study: stellar masses, SFRs, and quenched fractions. We emphasize that both the thermal and hybrid models were calibrated to reproduce the $z=0$ GSMF from Driver et al. (2022) and the $z\approx 0$ stellar mass – size relation from Hardwick et al. (2022) in the stellar mass range $10^{9}<M_{*}/\mathrm{M}_{\odot}<10^{11.3}$, as well as the $z\approx 0$ masses of SMBHs in massive galaxies. This implies that the properties of galaxies other than their $z=0$ stellar masses, half-mass radii, and SMBH masses should not necessarily be expected to match, especially at higher redshifts and outside the calibrated stellar mass range, where no observational constraints were applied to tune the simulations. Additionally, although the primary difference between the thermal and hybrid models lies in the AGN feedback prescription, we note that one of the SN feedback parameters also differs slightly. Namely, at all resolutions, the normalization of the gas density-dependent heating temperature $\Delta T_{\rm SN}(\rho_{\rm gas})\propto\rho_{\rm gas}^{2/3}$ is increased in the hybrid model, resulting in lower heating temperatures at fixed gas density and therefore weaker SN feedback. This adjustment was introduced to counterbalance the stronger AGN feedback in the hybrid model in intermediate-mass galaxies (see Huško et al. 2026 for details).
@@
  • At $z=0.1$, the GSMFs of the thermal and hybrid models again become indistinguishable, while the normalization of the SFMS is about $0.1$ dex lower in the hybrid model. Interestingly, the trend in quenched fractions reverses: the hybrid model now exhibits a higher quenched fraction, reaching $\approx 35$ per cent at $10^{9}\lesssim M_{*}/\mathrm{M_{\odot}}\lesssim 10^{10}$, compared to $\approx 20$ per cent in the thermal model. This is a consequence of the stronger AGN feedback produced by the hybrid model, which extends into the dwarf galaxy regime (see Huško et al., 2026) and will be studied in detail in future work. At higher stellar masses ($M_{*}\gtrsim 10^{11}~\mathrm{M_{\odot}}$), the quenched fractions in both models converge, with the hybrid model predicting marginally higher values. At even greater masses ($M_{*}\gtrsim 10^{11.5}~\mathrm{M_{\odot}}$), the quenched fractions in both models decline to around 50 per cent, suggesting that neither thermal energy injection nor kinetic jets can fully offset the strong cooling flows in the most massive galaxies. We note, however, that the quenched fractions at the high-mass end are particularly sensitive to uncertainties in stellar mass due to Eddington bias, which may artificially elevate their values (see Appendix D).
The hybrid model exhibits significantly lower quenched fractions at high redshifts due to weaker AGN feedback, resulting mainly from slower BH growth at high $z$, as well as from the fact that jets in the hybrid model require longer time-scales to quench the host galaxy (see fig. G1 in Chandro-Gómez et al. 2025). Although both the thermal and hybrid AGN feedback models use the modified Bondi-Hoyle-Lyttleton formula (with turbulence and vorticity corrections from Krumholz et al. 2006) to compute BH mass accretion rates, the latter further modifies the accretion rate by multiplying it by the accretion efficiency, a dimensionless parameter $0\leq\varepsilon_{\rm accr}\leq 1$. Physically, this parameter reduces the accretion rate into the BH event horizon relative to that at the scales of the BH accretion disc to account for mass loss due to accretion disc winds (see Huško et al., 2026, for details). At very high accretion rates (Eddington fractions greater than 1), which are readily reached in the colibre simulations at high redshift, $\varepsilon_{\rm accr}$ is set to 1 per cent, which slows down the BH mass growth in the hybrid model relative to the thermal model (see Huško et al., in preparation, for further details).
Over time, the lower quenched fractions in the hybrid model, combined with slightly weaker SN feedback, allow more stellar mass to form – resulting in a marginally elevated GSMF in the hybrid model at $M_{*}\gtrsim 10^{9.5}~\mathrm{M_{\odot}}$ by $z=2$. By $z=0.1$, the GSMFs in both models converge, as they were calibrated to match the same observational constraints. Mathematically, the lowz-z quenching in the hybrid model is expected to be stronger than in the thermal model, to ensure that the same amount of stellar mass is formed by $z=0$. Physically, this stronger quenching arises from the hybrid model’s more efficient AGN feedback, which at low redshifts is dominated by the kinetic jet mode, which experiences less radiative losses compared to thermal AGN feedback.
@@
Both models form similar amounts of stellar mass at $z\gtrsim 7$. Between $z=7$ and $z=2$, the CSMDs in the two models diverge, with the differences peaking at $z\approx 2$, where the hybrid model has formed $\approx 25$ per cent more stellar mass. From $z=2$ to 0, the CSMDs converge, with the differences approaching zero by $z\approx 0$, as both models were calibrated to reproduce the same observed $z=0$ GSMF.